OpenStudy (anonymous):

i got 9 but i think i did something wrong.. someone help plz. Determine the diameter of a circle that has an area of 18 cm2. If necessary round answers to nearest tenth

OpenStudy (btaylor):

A = πr^2 18 = πr^2 r^2 = 18/π \[r = \sqrt{\frac{18}{\pi}}\]

OpenStudy (anonymous):

2.39 ?

OpenStudy (anonymous):

or 5.7 ? @BTaylor

OpenStudy (btaylor):

it is 2.39